# Invariant derivations and trace bounds

### v3

@@ 5. Derivations on $T/W$ versus $W$-invariant derivations on $T$
Since $G^{\prime}$ is simply connected, $T/W$ is the product of an affine space and a torus (Prop. 3.4), and thus we may identify $T/W$ with an open subscheme of $\mathbb{A}^{r}$ where $r=\dim(T)$. Since open embeddings are étale, the sheaf of holomorphic differentials on $T/W$ is the pullback of the sheaf of holomorphic differentials on $\mathbb{A}^{r}$. Since the latter is free (with basis $df_{1},\ldots,df_{r}$ where $f_{1},\dots,f_{r}$ are the coordinate functions), so is $\Omega_{\mathbb{C}[T/W]}$. ∎
@@
and thus is determined by the symmetric matrix whose $(i,j)$-entry is $M_{A}(f_{i},f_{j})$. If we also have an isomorphism $(z_{1},\dots,z_{n})\colon T\xrightarrow{\sim}\mathbb{G}_{\mathrm{m}}^{n}$, then this matrix factors via the chain rule:
@@
is the $(k,l)$-entry of the symmetric matrix corresponding to the quadratic form $A$ in the basis corresponding to $\vec{z}$. In other words, one has
(5.7) $M_{A}(f_{i},f_{j})=(J^{\mathsf{T}}AJ)_{ij}$
where $J$ is the Jacobian matrix with entries $z_{i}\partial_{z_{i}}f_{j}$. After fixing a choice of $A$ and coordinates $f_{1},\ldots,f_{n}$, we will write simply $M$ for the matrix whose $(i,j)$-entry is given by (5.7).
@@
Suppose $G$ is simple (and simply connected). There is a unique Weyl-invariant symmetric bilinear form $b$ on the root lattice $Q$ such that $b(\alpha,\alpha)=2$ for each short root alpha; it is positive definite and integer-valued. Put $B$ for the matrix of $b$ with respect to some framing of $G$ with respect to $T$. Since $G$ is simply connected, the space $\mathfrak{t}^{*}$ is identified with the complexification $P\otimes\mathbb{C}$ of the weight lattice $P$ and one can take $f_{1},\ldots,f_{n}$ to be the fundamental characters. We take $A$ to be some positive integer multiple of $B$; any such choice will determine a real inner product on $P\otimes\mathbb{R}$. In the the Magma code for this paper and in the $G_{2}$, $F_{4}$, and $E_{8}$ examples in this part of the paper, we use $A:=|P/Q|\,B$. (Magma’s CoxeterForm, when passed a root datum with IsogenyClass := "SC", returns a matrix for $\frac{1}{2}B$ with respect to a basis consisting of fundamental weights.) Choosing a different scalar factor will scale the matrix $A$, but will leave the critical point computations otherwise unchanged, see Corollaries 5.11 and 5.12.
@@
$M=J^{T}J=\left(\begin{smallmatrix}z_{1}^{2}+z_{2}^{2}&z_{1}^{2}z_{2}+z_{1}z_{2}^{2}\\ z_{1}^{2}z_{2}+z_{1}z_{2}^{2}&2z_{1}^{2}z_{2}^{2}\end{smallmatrix}\right)=\left(\begin{smallmatrix}f_{1}^{2}-2f_{2}&f_{1}f_{2}\\ f_{1}f_{2}&2f_{2}^{2}\end{smallmatrix}\right),$
@@
Since $M$ is symmetric, it suffices to consider only the claim involving a column. The $i$-th column vanishes if and only if $D_{j}(f_{i})(t)=0$ for all $j$; since we have seen that the $D_{j}$ form a basis of the submodule of $\operatorname{Der}(\mathbb{C}[T/W])$ corresponding to $W$-invariant derivations, this is equivalent to having $D(f_{i})(t)=0$ for every $W$-invariant derivation $D$ on $T$. This is equivalent by Prop. 4.2 to $t$ being a critical point for $f_{i}|_{T}$, which in turn is equivalent to being a critical point for $f_{i}$ on $G$ (Lemma 4.1). ∎
@@
Let $G$ be a semisimple simply connected algebraic group over $\mathbb{C}$. The corners are the images of those $t\in T(\mathbb{C})$ that are critical points for every $G$-invariant function.
@@
Although one no longer has an application to minimization or maximization, the module of invariant derivations is still a natural object to consider in finite characteristic. There are two places in which the above construction may fail. For one, it may not be possible to choose a nonsingular invariant form $A$. In that case, $\det(A)=0$, so $\det(M)=0$, and the rows of $M$ do not even generate the space of meromorphic invariant derivations.
@@
This leads to a natural question: Under what circumstances is the module of invariant derivations still free over the ring of invariants in general? We computed spanning sets [^2] See the Magma program bad-char in the git repository. for the modules of invariant derivations for $\operatorname{SL}_{5}$ in characteristic 5 and $\operatorname{SL}_{6}$ in characteristics 2 and 3; in each case have found that the module is not free at the identity. It thus seems likely that the module is almost never free in bad characteristic, with only sporadic exceptions (e.g., $\operatorname{SL}_{2}$ and $\operatorname{SL}_{4}$ in characteristic 2 and $\operatorname{SL}_{3}$ in characteristic 3).
@@
There is an alternate approach to construct derivations. There is another natural way to associate an operator $C_{A}$ on $\mathbb{C}[T]^{W}\cong\mathbb{C}[G]^{G}$ to $A$, in such a way as to be diagonal on the irreducible characters.
This operator is not quite $\mathcal{D}_{A}$, but is closely related. For $G$ simply connected, let $\delta$ be the Weyl denominator
$\delta=z^{\rho}\prod_{\alpha\in\Phi^{+}}(1-z^{-\alpha}).$
(See [Bou02, §VI.3.3] for general results on $\delta$.) By the Weyl character formula, multiplying a character by $\delta$ gives the antisymmetrization of a monomial, and thus gives an eigenfunction for $\mathcal{D}_{A}$. In other words, $\delta^{-1}\mathcal{D}_{A}\delta$ is diagonal on the irreducible characters, suggesting (after subtracting off the eigenvalue at $1$) the definition
$C_{A}:=\delta^{-1}\mathcal{D}_{A}\delta-\delta^{-1}\mathcal{D}_{A}(\delta).$
Since $\mathcal{D}_{A}(1)=0$, we find that in fact
@@
This has the disadvantage of forcing us to leave $\mathbb{C}[T]^{W}$, but this can be avoided: since $M_{A}$ is a derivation in the first argument, we have
$\delta^{-k}M_{A}(\delta^{k},{-})=k\delta^{-1}M_{A}(\delta,{-})$
for any integer $k$, and thus in particular may also write
(5.14) $C_{A}=\mathcal{D}_{A}+\frac{1}{2}\Delta^{-1}M_{A}(\Delta,{-}),$
where
$\Delta:=\delta^{2}=z^{2\rho}\prod_{\alpha\in\Pi^{+}}(1-z^{-\alpha})^{2}=\prod_{\alpha\in\Pi^{+}}(z^{\alpha}-2+z^{-\alpha}).$
The latter expression gives a well-defined element of $\mathbb{C}[G]^{G}$ for any reductive $G$, and thus we could alternatively define $C_{A}$ via (5.14) for any reductive $G$.
@@
The operator $C_{A}$ is a second-order operator on $\mathbb{C}[T]^{W}$ with associated biderivation $M_{A}$. Moreover, it is diagonal in the basis of irreducible characters, with eigenvalue $A(\lambda+\rho)-A(\rho)$ on the character with highest weight $\lambda$.
@@
We first show that it is diagonal on the irreducible characters. If $\rho$ is a weight of $G$, then we have
$C_{A}=\delta^{-1}\mathcal{D}_{A}\delta-\delta^{-1}\mathcal{D}_{A}(\delta),$
and thus the claim follows from the Weyl character formula. Otherwise, there is a double cover $G^{+}$ for which $\rho$ is a weight, and (since isogenies are étale) an induced pairing $A^{+}$. Any irreducible character of $G^{+}$ is an irreducible character for $G$, and the action of $C_{A^{+}}$ agrees with the action of $C_{A}$ where both are defined.
We have thus seen that $C_{A}$ is a holomorphic operator. That it is a second-order operator with associated biderivation $M_{A}$ then follows from the fact that
$C_{A}-\mathcal{D}_{A}=\frac{1}{2}\Delta^{-1}M_{A}(\Delta,{-})$
is a derivation. ∎
@@
If $G$ has maximal compact subgroup $K$ and $T$ contains a maximal torus $T_{K}$ of $K$, then Haar measure on $K$ induces a $W$-measure on $T_{K}$, and the density of that measure relative to Haar measure on $T_{K}$ is (up to an irrelevant scalar multiple) nothing other than $\Delta$. In particular, since $\mathcal{D}_{A}$ is self-adjoint relative to Haar measure on $T_{K}$, $C_{A}$ is self-adjoint relative to the measure induced from $K$, which is why its eigenfunctions form an orthonormal basis relative to that measure.
@@
Note that both $\mathcal{D}_{A}$ and $C_{A}$ are reasonably straightforward to compute on the character ring: $\mathcal{D}_{A}$ is diagonal on monomials and $C_{A}$ is diagonal on irreducible characters. So to apply the operators to an element of the character ring, it suffices to expand it into irreducibles (or monomials) and apply the relevant diagonal operator. This gives the coefficients of the matrix $M$ as an expansion in irreducibles (or monomials), and thus one reduces to the problem of expressing those coefficients as polynomials in the fundamental characters.
@@
Consider the case where $G$ is compact of type $G_{2}$, where the roots are numbered as in the diagram
@@
The critical points of $f_{2}$, the character of the adjoint representation, is the locus cut out by the second row (or column) of $M$, which has four components, three of which are the corners. The remaining component is $(f_{1},f_{2})=(7/9,10/27)$. Since no critical point has $f_{2}$ less than $-2$ and $-2$ occurs as a value of $\operatorname{Tr}\operatorname{Ad}$ on the compact group (because $-1$ belongs to the Weyl group, Example 3.1), we conclude that $\min\operatorname{Tr}\operatorname{Ad}=-2$, verifying Theorem 2.3 in this case.
The critical points of $f_{1}$, the character of the highest short root representation, is the locus cut out by the first row (or column) of $M$. This consists solely of the corners (5.19).
@@ 6. Critical points from the compact group
$M^{\sigma}_{A}(g,f)=\sigma(M^{\sigma}_{A}(f,g)).$
@@
Continue Example 5.8, where $G$ is simple and simply connected. On the root system, $-1$ and $-w_{0}$ for $w_{0}$ the longest element of the Weyl group $W$ are in the same coset of $W$. Since $-w_{0}$ permutes the simple roots, $\sigma$ permutes the fundamental characters in the same way, and $M^{\sigma}$ is obtained from $M$ by permuting the rows of $M$ in the same manner.
@@
Let $V\cong\mathbb{R}^{n}$ be Euclidean space, and let $\Gamma$ be a discrete subgroup of the “(real) affine orthogonal group” $V\rtimes O(V)$ with finite image in $O(V)$. The action of $\Gamma$ on $V$ extends naturally to an action on $V\otimes\mathbb{C}$, and the quotient $X:=(V\otimes\mathbb{C})/\Gamma$ has a natural structure as an affine scheme over $\mathbb{R}$. Indeed, the quotient by the lattice $\Gamma\cap V$ is a product of an affine space and an algebraic torus, with real structure such that the real locus is the product of a vector space and a topological torus. This is clearly an affine scheme, and $(V\otimes\mathbb{C})/\Gamma$ is the quotient of this affine $\mathbb{R}$-scheme by the real and finite group $\Gamma/(\Gamma\cap V)$.
@@
Let $T$ be a maximal torus of $G$ containing $g$ and write $W$ for $N_{G}(T)/T$. Then by Steinberg [Ste65, 6.4] $\operatorname{Class}(G)\cong T/W$ over $\mathbb{R}$, so the hypothesis on $g$ implies that the image of $g$ in $(T/W)(\mathbb{C})$ belongs to $(T/W)(\mathbb{R})$. The theorem gives: $M^{\sigma}(g)$ is negative semidefinite if and only if there exists a $t\in T(\mathbb{R})$ with the same image in $(T/W)(\mathbb{R})$. Applying Steinberg again, we see that this condition is equivalent to $g$ being $G(\mathbb{C})$-conjugate to an element of $T(\mathbb{R})$. ∎

### v4

@@ 5. Derivations on $T/W$ versus $W$-invariant derivations on $T$
Since $G^{\prime}$ is simply connected, by Prop. 3.4bbii we may identify $T/W$ with an open subscheme of $\mathbb{A}^{r}$ where $r=\dim T$. Since open embeddings are étale, the sheaf of Kähler differentials on $T/W$ is the pullback of the sheaf of Kähler differentials on $\mathbb{A}^{r}$. Since the latter is free (with basis $df_{1},\ldots,df_{r}$ where $f_{1},\dots,f_{r}$ are the coordinate functions), so is $\Omega_{\mathbb{C}[T/W]}$. ∎
@@
and thus is determined by the symmetric matrix $M$ whose $(i,j)$-entry is $M_{A}(f_{i},f_{j})$. If we also have an isomorphism $(z_{1},\dots,z_{n})\colon T\xrightarrow{\sim}\mathbb{G}_{\mathrm{m}}^{n}$, then this matrix factors via the chain rule:
@@
is the $(k,l)$-entry of the symmetric matrix corresponding to the quadratic form $A$ in the basis corresponding to $\vec{z}$. In other words, the matrix $M$ (which depends on the choice of $A$ and the coordinates $f_{1},\ldots,f_{n}$) is
$M=J^{\mathsf{T}}AJ,$
where $J$ is the Jacobian matrix with $(i,j)$-entry $z_{i}\partial_{z_{i}}f_{j}$.
@@
Suppose $G$ is simple (and simply connected). There is a unique Weyl-invariant symmetric bilinear form $b$ on the root lattice $Q$ such that $b(\alpha,\alpha)=2$ for each short root alpha; it is positive definite and integer-valued. Put $B$ for the matrix of $b$ with respect to some pinning of $G$ with respect to $T$. Since $G$ is simply connected, the space $\mathfrak{t}^{*}$ is identified with the complexification $P\otimes\mathbb{C}$ of the weight lattice $P$ and one can take $f_{1},\ldots,f_{n}$ to be the fundamental characters. We take $A$ to be some positive integer multiple of $B$; any such choice will determine a real inner product on $P\otimes\mathbb{R}$. In the the Magma code for this paper and in the $G_{2}$, $F_{4}$, and $E_{8}$ examples in this part of the paper, we use $A:=|P/Q|\,B$. (Magma’s CoxeterForm, when passed a root datum with IsogenyClass := "SC", returns a matrix for $\frac{1}{2}B$ with respect to a basis consisting of fundamental weights.) Choosing a different scalar factor will scale the matrix $A$, but will leave the critical point computations otherwise unchanged, see Corollaries 5.10 and 5.12.
@@
$M=J^{\mathsf{T}}J=\left(\begin{smallmatrix}z_{1}^{2}+z_{2}^{2}&z_{1}^{2}z_{2}+z_{1}z_{2}^{2}\\ z_{1}^{2}z_{2}+z_{1}z_{2}^{2}&2z_{1}^{2}z_{2}^{2}\end{smallmatrix}\right)=\left(\begin{smallmatrix}f_{1}^{2}-2f_{2}&f_{1}f_{2}\\ f_{1}f_{2}&2f_{2}^{2}\end{smallmatrix}\right),$
@@
For semisimple $g\in G(\mathbb{C})$, it makes sense to write $M(g)$ for the matrix with complex entries obtained by evaluating each entry of $M$ (which belongs to $\mathbb{C}[T/W]$) at the image of $g$ in $(T/W)(\mathbb{C})$.
@@
Since $M$ is symmetric, it suffices to consider only the claim involving a column. The $i$-th column vanishes if and only if $D_{j}(f_{i})(t)=0$ for all $j$; since we have seen that the $D_{j}$ form a basis of the submodule of $\operatorname{Der}\mathbb{C}[T/W]$ corresponding to $W$-invariant derivations, this is equivalent to having $D(f_{i})(t)=0$ for every $W$-invariant derivation $D$ on $T$. This is equivalent by Prop. 4.2 to $t$ being a critical point for $f_{i}|_{T}$, which in turn is equivalent to being a critical point for $f_{i}$ on $G$ (Lemma 4.1). ∎
Definition 5.11.
@@
Since $(T/W)(\mathbb{C})$ is in bijection with conjugacy classes of semisimple elements of $G(\mathbb{C})$, one could equally well refer to the conjugacy classes corresponding to corners as corners themselves.
@@
Let $G$ be a semisimple simply connected algebraic group. The corners are the images of those $t\in T(\mathbb{C})$ that are critical points for every $G$-invariant function.
@@
Although one no longer has an application to minimization or maximization, the module of invariant derivations is still a natural object to consider over a field of prime characteristic. There are two places in which the above construction may fail. For one, it may not be possible to choose a nonsingular invariant form $A$. In that case, $\det(A)=0$, so $\det(M)=0$, and the rows of $M$ do not even generate the space of rational invariant derivations.
@@
This leads to a natural question: Under what circumstances is the module of invariant derivations still free over the ring of invariants in general? We computed spanning sets [^2] See the Magma program bad-char in the git repository. for the modules of invariant derivations for $\operatorname{SL}_{5}$ in characteristic 5 and $\operatorname{SL}_{6}$ in characteristics 2 and 3; in each case we have found that the module is not free at the identity. It thus seems likely that the module is rarely free when the characteristic is not “very good”, with only sporadic exceptions (e.g., $\operatorname{SL}_{2}$ and $\operatorname{SL}_{4}$ in characteristic 2 and $\operatorname{SL}_{3}$ in characteristic 3).
@@
There is an alternate approach to construct derivations. There is another natural way to associate an operator $C_{A}$ on $\mathbb{C}[T]^{W}\cong\mathbb{C}[G]^{G}$ to $A$, in such a way as to be diagonal on the characters of irreducible representations. Define
(5.14) $C_{A}:=\mathcal{D}_{A}+\frac{1}{2}\Delta^{-1}M_{A}(\Delta,{-})$
where
$\Delta:=\prod_{\alpha\in\Phi^{+}}(e^{\alpha}-2+e^{-\alpha})\quad\in\mathbb{C}[T]^{W}.$
Here we are writing an unadorned symbol such as $\alpha$ for roots and weights (viewed as elements of $\mathfrak{t}^{*}$, which is canonically isomorphic to $\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})\otimes\mathbb{C})$ and $e^{\alpha}$ for the corresponding element of $\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})$.
We now describe $C_{A}$ in a special case. Recall the Weyl denominator
$\delta:=e^{\rho}\prod_{\alpha\in\Phi^{+}}(1-e^{-\alpha})$
where $\rho$ is half the sum of the positive roots. If $\rho$ corresponds to an element of $\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})$, then $\delta$ belongs to $\mathbb{C}[T]$ and is anti-invariant under the Weyl group [Bou02, §VI.3.3, Prop. 2(i)], and moreover we have
$\delta^{2}=e^{2\rho}\prod_{\alpha\in\Phi^{+}}(1-e^{-\alpha})^{2}=\Delta.$
Lemma 5.15.
If the half sum of the positive roots, $\rho$, corresponds to an element of $\operatorname{Hom}(T,\mathbb{G}_{\mathrm{m}})$, then
(5.16) $C_{A}=\delta^{-1}\mathcal{D}_{A}\delta-\delta^{-1}\mathcal{D}_{A}(\delta).$
Proof.
Since $M_{A}$ is a derivation in the first argument, we have
$\delta^{-k}M_{A}(\delta^{k},{-})=k\delta^{-1}M_{A}(\delta,{-})$
for any integer $k$. Combining this with the definition (5.14) of $C_{A}$ gives
@@
Since $\mathcal{D}_{A}$ is a homogeneous quadratic polynomial in derivations (which each kill 1), we have $\mathcal{D}_{A}(1)=0$. Combining this with the expression for $M_{A}$ as a biderivation arising from $\mathcal{D}_{A}$ as in Lemma 5.3b, we find that $C_{A}$ equals (5.16). ∎
The inspiration for the definition (5.14) came from (5.16). By the Weyl character formula, multiplying an irreducible character by $\delta$ gives the antisymmetrization of a monomial, and thus gives an eigenfunction for $\mathcal{D}_{A}$. In other words, $\delta^{-1}\mathcal{D}_{A}\delta$ is diagonal on the irreducible characters, which suggested (after subtracting off the eigenvalue at $1$) (5.16).
@@
The operator $C_{A}$ is a second-order differential operator on $\mathbb{C}[T]^{W}$ with associated biderivation $M_{A}$. Moreover, it is diagonal in the basis of irreducible characters, with eigenvalue $A(\lambda+\rho)-A(\rho)$ on the character with highest weight $\lambda$.
@@
We first show that it is diagonal on the irreducible characters. If $\rho$ is a character of $G$, then $C_{A}$ is given by (5.16) and the claim follows from the Weyl character formula. Otherwise, there is a double cover $G^{+}$ for which $\rho$ is a weight, and (since isogenies are étale) a corresponding induced pairing $A^{+}$. Any irreducible character of $G$ is an irreducible character for $G^{+}$, and the action of $C_{A^{+}}$ agrees with the action of $C_{A}$ where both are defined.
We have thus seen that $C_{A}$ preserves $\mathbb{C}[T]^{W}$. Since $\mathcal{D}_{A}$ is a second-order differential operator, and $C_{A}-\mathcal{D}_{A}$ is a derivation by (5.14), $C_{A}$ is also a second-order differential operator with the same associated biderivation as $\mathcal{D}_{A}$. ∎
@@
If $G$ has maximal compact subgroup $K$ and $T$ contains a maximal torus $T_{K}$ of $K$, then Haar measure on $K$ induces a $W$-invariant measure on $T_{K}$, and the density of that measure relative to Haar measure on $T_{K}$ is (up to an irrelevant scalar multiple) nothing other than $\Delta$. In particular, since $\mathcal{D}_{A}$ is self-adjoint relative to Haar measure on $T_{K}$, $C_{A}$ is self-adjoint relative to the measure induced from $K$, which is why its eigenfunctions form an orthonormal basis relative to that measure.
@@
Note that both $\mathcal{D}_{A}$ and $C_{A}$ are reasonably straightforward to compute on $\mathbb{C}[G]^{G}$: $\mathcal{D}_{A}$ is diagonal on monomials and $C_{A}$ is diagonal on irreducible characters. So to apply the operators to an element of $\mathbb{C}[G]^{G}$, it suffices to expand it into irreducibles (or monomials) and apply the relevant diagonal operator. This gives the coefficients of the matrix $M$ as an expansion in irreducibles (or monomials), and thus one reduces to the problem of expressing those coefficients as polynomials in the fundamental characters.
@@
Consider the case where $G$ is simple of type $G_{2}$, where the roots are numbered as in the diagram
@@
The critical points of $f_{2}$, the character of the adjoint representation, are the locus cut out by the second row (or column) of $M$, which has four components, three of which are the corners. The remaining component is $(f_{1},f_{2})=(7/9,10/27)$. Since no critical point has $f_{2}$ less than $-2$ and $-2$ occurs as a value of $\operatorname{Tr}\operatorname{Ad}$ on the compact group (because $-1$ belongs to the Weyl group, Example 3.1), we conclude that $\min\operatorname{Tr}\operatorname{Ad}=-2$, verifying Theorem 2.3 in this case.
The critical points of $f_{1}$, the character of the highest short root representation, are the locus cut out by the first row (or column) of $M$. This consists solely of the corners (5.21).
@@ 6. Critical points from the compact group
$M^{\sigma}_{A}(g,f)=\sigma(M^{\sigma}_{A}(f,g))$
for all $f,g\in\mathbb{C}[T/W]$.
@@
Continue Example 5.7, where $G$ is simple and simply connected. Acting on the root system, $-1$ and $-w_{0}$ for $w_{0}$ the longest element of the Weyl group $W$ are in the same coset of $W$. Since $-w_{0}$ permutes the simple roots, $\sigma$ permutes the fundamental characters in the same way, and $M^{\sigma}$ is obtained from $M$ by permuting the rows of $M$ in the same manner.
@@
Let $V\cong\mathbb{R}^{n}$ be Euclidean space, and let $\Gamma$ be a discrete subgroup of the “(real) affine orthogonal group” $V\rtimes O(V)$ with finite image in $O(V)$. The action of $\Gamma$ on $V$ extends naturally to an action on $V\otimes\mathbb{C}$, and the quotient $X:=(V\otimes\mathbb{C})/\Gamma$ has a natural structure as an affine scheme over $\mathbb{R}$ where the affine space is associated to the orthogonal complement of $\mathbb{R}(\Gamma\cap V)$. Indeed, the quotient by the lattice $\Gamma\cap V$ is a product of an affine space and an algebraic torus, with real structure such that the real locus is the product of a vector space and a topological torus. This is clearly an affine scheme, and $(V\otimes\mathbb{C})/\Gamma$ is the quotient of this affine $\mathbb{R}$-scheme by the real and finite group $\Gamma/(\Gamma\cap V)$.
@@
Remark.
Lemma 6.9 is reminiscent of the main theorem 0.10 in [PS85], but that result concerns a linear representation of a compact group rather than an affine representation of a discrete group. When our group is finite, the lemma is a special case of their result, with a very similar proof. (They not only allow the group to be compact, but give the appropriate condition at singular points of the quotient.) Presumably there is a common generalization to affine representations of (topologically) closed subgroups of $V\ltimes O(V)$ with closed image in $O(V)$.
@@
Let $T$ be a maximal torus of $G$ containing $g$ and write $W$ for $N_{G}(T)/T$. Then by Steinberg [Ste65, 6.4] $\operatorname{Class}(G)\cong T/W$ over $\mathbb{R}$, so the hypothesis on $g$ implies that the image of $g$ in $(T/W)(\mathbb{C})$ belongs to $(T/W)(\mathbb{R})$. Theorem 6.7 gives: $M^{\sigma}(g)$ is negative semidefinite if and only if there exists a $t\in T(\mathbb{R})$ with the same image in $(T/W)(\mathbb{R})$. Applying Steinberg again, we see that this condition is equivalent to $g$ being $G(\mathbb{C})$-conjugate to an element of $T(\mathbb{R})$. ∎